OpenStudy (anonymous):

Due tomarrow! 1.car has a mass of 1.72×103kg.If the force acting on the car is 6.44×103N to the east,what is the car’s acceleration?Answer units of m/s2 3.What is the acceleration of a 56 kg block of cement when pulled sideways with a net force of 526 N?Answer in units of m/s2 4.Sprinting near the end of a race,a runner with a mass 62 kg accelerates from a speed of 12m/s to a speed of 13m/s in 31s.To gain speed the runner produces a backward force on the ground,so that the ground pushes the runner forward,providing the force necessary for the acceleration.Calcullate average force.

OpenStudy (deepolisnoob):

1.) \[F=ma\] Your \[F=6.44*10^{3}N\] and your \[m=1.72*10^{3}kg\] You plug those values into you equation, \[ 6,440N=1,720kg*a\] I simplified the scientific notation. Get your a by itself, so divide both sides by 1,720kg. \[\frac{ 6,440N }{ 1,720kg }=a\] \[3.74m/s^{2}=a\] 3.) \[F=ma\] \[526N=56kg*a\] \[\frac{526N}{56kg}=a\] \[a=5.39m/s^{2}\] 4.) \[F=ma\] \[a=\frac{\Delta v}{\Delta t}\] \[a=\frac{13m/s-12m/s}{31s}\] \[a=\frac{1m/s}{31s}\] \[a=0.032m/s^{2}\] \[F=62kg*0.032m/s^{2}\] \[F=1.984N\]

OpenStudy (anonymous):

it say # 1 was wrong and # 3 was too and the answer to 3 is figured out the same was but the answer is 9.39285